OpenStudy (ashontae19):

CAN SOMEONE PLEASE SHOW ME THE STEPS I WILL MEDAL let f(x)=-2x+4 and g(x)=-6x-7. find f(x)-g(x)

OpenStudy (ashontae19):

@mathmale

OpenStudy (ashontae19):

@Michele_Laino

OpenStudy (michele_laino):

it is simple, by definition, we can write: \[\Large \begin{gathered} f\left( x \right) - g\left( x \right) = - 2x + 4 - \left( { - 6x - 7} \right) = \hfill \\ \hfill \\ = - 2x + 4 + 6x + 7 = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

@ashontae19

OpenStudy (ashontae19):

would it be 4x+11?

OpenStudy (michele_laino):

that's right!

OpenStudy (ashontae19):

ok thanxs so much

OpenStudy (michele_laino):

:)